OpenStudy (anonymous):

if f(X)=1/9x-2, what isf-1(x)?

OpenStudy (anonymous):

The way I would solve this problem would be replacing f(x) with y. This would give you y=1/(9x -2) Then you would switch y's and x's giving you x=1/(9y-2). Now if you solve for the new y, it will give you f-1(x). I will solve it out for you. x(9y-2)=1 9y-2=1/x 9y=1/x +2 Answer: y=1/(9x)+2/9 or f-1(x)=1/(9x)+2/9 Note: I am not sure if you intended f(x) to be 1/(9x-2), 1/(9x) -2